# PIPA: A Unified Evaluation Protocol for Diagnosing Interactive Planning Agents

### v1

@@ 3.1.5 Task Completion ($\mathcal{R}$)
The primary objective of a task agent is to effectively accomplish goal-oriented tasks. The task completion metric adheres to the task-specific performance criteria established by each benchmark, since benchmark scenarios typically define their own evaluation frameworks to determine what constitutes a “completed task.” Specifically, most metrics in diverse benchmarks demonstrated in Table 5 can be regarded as task completion metrics since they mainly focus on whether agents successfully accomplish the given task.
@@ 3.1.6 Agent Interaction Pattern
While prior work has typically reported on dataset-level statistics (e.g., basic information in Table 1), it has rarely addressed how agents actually interact within their environments. We suggest reporting the following information, providing an overview of the interactive tendency of employed agents in their specific scenarios.
The Number of Turns
Different from offline dialogue datasets where a fixed number of turns is provided, recently employed online evaluation scenarios typically omit details about the agent interactivity. However, the number of interactive turns serves as a critical indicator of the conversational or interactive dynamics. In general, planning scenarios with more turns tend to require further management of user constraints, suggesting the need for sustained engagement and state tracking.
The Number of Steps
Steps refer to internal actions or reasoning hops, such as goal decomposition or tool use. This feature reflects the agent’s ability to handle complex tasks that require breaking down overarching goals into smaller, manageable actions. Fewer steps may suggest shallow or overly generic reasoning, while more steps indicate finer control. Correlating the count of steps with performance and the length of interaction can reveal whether agents are making meaningful progress or getting stuck in unproductive loops.
@@ 4.1 Benchmarks
# of Dialogues # of Tools # of Database
TravelPlanner (2024a) 180 7 3,865,195 total, 3,827,361 max for a tool
$\tau$-Bench-Airline (2025) 50 13 500 users, 300 flights, 2,000 reservations
$\tau$-Bench-Retail (2025) 115 15 500 users, 50 products, 1,000 orders
To verify the effectiveness of the proposed evaluation protocol, we employ challenging task planning benchmarks that conform to the prerequisite assumptions described in Section 3 (i.e., predefined tools and external database). Specifically, we utilize three task domains in two benchmarks: TravelPlanner (Xie et al., 2024a), a single-turn dataset in the itinerary domain, and $\tau$-Bench (Yao et al., 2025), a multi-turn dataset encompassing the airline and retail domains. Table 1 presents the statistics of these benchmarks. For the evaluation metric for task completion ($\mathcal{R}$), pass rate is used for TravelPlanner, and pass^k is used in $\tau$-Bench. Refer to Table 5 for the description of each metric.
4.2 LLMs for Agent, User Simulator, and Performance Evaluator
@@
[…]
@@
Table 2: PIPA evaluation result. The relative performance ranking among agents for each metric is differentiated with colors (i.e., darker color indicates more competitive performance; different colors are applied for proprietary and open-source models). Best average performance is indicated in bold, and second-best is underlined.
@@
Avg. Turns Avg. Steps Avg. Turns Avg. Steps Avg. Turns Avg. Steps
@@
Open-Source Models
@@
We tested diverse task planning agents consisting of five proprietary models (gpt-4o, gpt-4o-mini, gpt-3.5-turbo, gemini-1.5-flash, sonnet-3.5) and four open-source models (mistral-large 444123B-sized model: https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, mixtral-8x7B, llama-3.3-70B, qwen2.5-72B). Since $\tau$-Bench requires multi-turn interactions between a user and agent, we employ gpt-4o as a user simulator with the same instruction as the original work (Yao et al., 2025). For the LLM evaluator leveraged in measuring PIPA metrics, we employ llama-3.3-70B (Grattafiori et al., 2024) [^5] We confirm llama-3.3-70B shows better performance compared to gpt-4o-mini through our preliminary investigation.. All prompts used in our experiments are listed in Appendix B.
@@ 4.3 Results
Behavioral Stage Diagnosis
Table 2 presents the evaluation results following PIPA  [^6] We report agents’ erroneous behaviors at each stage in Appendix C.. We differentiate the color ranking of proprietary and open-source models to observe diverse patterns. Along with individual metrics in PIPA, we report the average score to see the overall behavioral performance.
A notable observation is that the conventional task completion metric ($\mathcal{R}$) does not necessarily reflect the performance of intermediate phases. For instance, while qwen2.5-72B achieves the highest performance on task completion (.38) among open-source models in $\tau$-Bench-Retail, its PIPA metric results lag behind other models. Moreover, although most agents in TravelPlanner show the same zero task completion performance, each model has different intermediate capability patterns according to PIPA metrics. These observed patterns imply a potential discrepancy between task completion performance values and qualitative user preferences, as user satisfaction is influenced not only by task completion but also by the quality of interactions across all phases. We introduce a human study for verifying this discrepancy in Section 5.1.
Interaction Pattern Diagnosis
We summarize each agent’s interaction pattern in Table 3. While TravelPlanner is characterized by single-turn interactions with numerous intermediate steps prior to producing a final response, $\tau$-Bench exhibits fewer intermediate steps per turn but involves multiple turns. These patterns reveal certain behavioral tendencies among the agents; for instance, competitive agents (e.g., sonnet-3.5, llama-3.3-70B) appear to engage in more extensive internal reasoning within individual turns, yet participate in fewer turns overall in multi-turn benchmarks. This observation suggests that, in adapting TravelPlanner for multi-turn scenarios, careful consideration must be given to balancing the number of turns and steps within each turn. Given that service providers often target different application scenarios, analyzing interaction patterns across benchmarks can inform the strategic selection of those most aligned with specific deployment contexts.
@@ 5 Analyses and Discussions
5.1 Relationship Between Task Completion
and User Satisfaction
Our hypothesis regarding the results presented in Table 2 is that factors beyond task completion may have significantly influenced overall user satisfaction. To investigate this, we conduct a human study involving 16 graduate-level participants across the following two controlled scenarios:
(1) Same $\mathcal{R}$ Different AVG.
We select gpt-4o-mini and mistral-large in $\tau$-Bench-Airline, as these two models exhibit identical task completion performance (.26), yet differ in their PIPA average scores (.64 and .56, respectively).
@@ 5.1 Relationship Between Task Completion
We select gemini-1.5-fsh. and llama-3.3-70B in $\tau$-Bench-Retail, as these two models exhibit identical PIPA average score (.64), yet differ in their task completion performance (.15 and .36, respectively).
Participants are presented with two multi-turn conversations generated by different agents interacting with a gpt-4o user simulator. They are asked to indicate which conversation they find more satisfactory [^7] Detailed instructions are provided in Section D.1, and we conduct qualitative analysis for participants’ decisions in Section D.2.. If both conversations are perceived as equally (un)satisfactory, participants are given the option to select “no preference.” Each participant evaluates a total of six conversation pairs (randomly sampled three pairs of conversations, respectively).
Figure 3: Human study results for measuring preferred interactions.
The results of the human evaluation are presented in Figure 3. For both scenarios, we conduct binomial tests by excluding “no preference” responses, yielding statistically significant results (p-value $<$ 0.05 in both cases). In the Airline domain study, we observe that participants tend to favor conversations with higher average score when task completion is held constant. Conversely, in the Retail domain study, task completion still remains a strong predictor of user satisfaction when the average score is identical. These findings highlight the complementary roles of mostly-used task completion and PIPA metrics in capturing different facets of user satisfaction, as demonstrated in Table 2. Our study thus supports the use of PIPA as a meaningful proxy for human preferences and motivates future research into evaluation frameworks that go beyond simple task outcomes.
5.2 Can Mixing Agents Lead to Better
Performance?
Building upon the insights from previous experiments, we are further motivated to explore the potential for enhancing performance and user satisfaction by leveraging diverse agents strengths within the agentic pipeline. To facilitate this investigation, we separate the agent’s interactive process into two distinct components: (1) intermediate understanding and (2) response generation. This separation enables targeted improvements, wherein state consistency and tool efficiency are addressed in the first stage, whereas observation alignment and policy alignment are the focus of the second. Specifically, we employ one agent to understand user requests and call proper tools, then replace it with the other when starting to generate responses.
[…]
@@ 5.2 Can Mixing Agents Lead to Better
For faster investigation of this experimental study, we utilize FP8-quantized versions of three open-source models (llama-3.3-70B, qwen2.5-72B, mistral-large), alongside llama-3.3-70B as the user simulator, and evaluate on $\tau$-Bench-Airline with five repeated runs per configuration. Selected results are presented in Table 4. While acknowledging that arbitrary combinations of agents do not consistently yield improved performance, even leading to drop in performance, certain pairings demonstrate notable enhancements in either task completion or the average PIPA score. Notably, combining qwen2.5-72B (for intermediate understanding) with llama-3.3-70B (for response generation) leads to a substantial improvement in task completion (.263 $\rightarrow$ .361) without compromising the average PIPA score too much (.628 $\rightarrow$ .624). These findings suggest a promising direction for future research to strategically arrange agents across different stages of the task planning pipeline. For additional details and comprehensive experimental results beyond the scope of this discussion, please refer to Appendix E.
@@ 6 Conclusion
We introduce PIPA, a domain- and task-agnostic evaluation protocol grounded in a partially observable Markov Decision Process (POMDP) framework, to capture how user satisfaction is shaped by both final outputs and intermediate behaviors in interactive task planning. Our experiments reveal that different models excel at various stages, underscoring the need for a holistic view of agent performance. Human studies also support that focusing solely on task completion can be misleading; stage-specific indicators often correlate more strongly with the user’s overall experience. These findings demonstrate that PIPA offers a principled, diagnostic lens to help researchers and practitioners design and benchmark more robust, transparent, and user-centric planning agents.
@@ Appendix A Evaluation Metrics Across Task Planning Benchmarks
MultiWOZ (2018) Success Rate The system answered all requested attributes.
@@
AgentBench (2024) Domain-Specific Different metrics such as Success Rate, Win Rate,
@@
WebLINX (2024) Intent Match Boolean indicator for correct intent prediction.
@@
FlowBench (2024) Tool Invocation Correctly identified tool configs (P/R/F1).
@@
$\tau$-Bench (2025) Pass@k Boolean success indicator for k attempts.
Pass^k Whether all k trials are successful.
@@
WebShop (2022) Task Score The average reward across episodes.
@@
TravelPlanner (2024a) Delivery Rate The agent completed tasks within limited steps.
@@ Appendix B Prompts
Basically, we keep the same agent prompt as the original paper (Xie et al., 2024a; Yao et al., 2025). The following prompts are evaluation prompts adopted in PIPA evaluation.
@@ 3. Accuracy and truthfulness
"state_id": "1",
"justification": "brief explanation...",
"score": "0"
"state_id": "2",
"justification": "brief explanation...",
"score": "1"
@@ 4. Consistency with previous details
"response_id": "1",
"justification": "brief explanation...",
"score": "0"
"response_id": "2",
"justification": "brief explanation...",
"score": "1"
@@
"policy_id": "1",
"justification": "Mention specific policy and turn message, if violated. Otherwise, mention all policies are adhered to.",
"score": "0"
"policy_id": "2",
"justification": "Mention specific policy and turn message, if violated. Otherwise, mention all policies are adhered to.",
"score": "1"

### v2

@@ 3.2.5 Task Completion ($\mathcal{R}$)
As in many agent studies, the primary objective of a task agent is to effectively accomplish goal-oriented tasks. The task completion metric adheres to the task-specific performance criteria established by each benchmark, since benchmark scenarios typically define their own evaluation frameworks to determine what constitutes a “completed task.” Specifically, most metrics in diverse benchmarks demonstrated in Table 5 can be regarded as task completion metrics since they mainly focus on whether agents successfully accomplish the given task. Again, it is important to note that AURA incorporates existing task completion metrics to complement the holistic evaluation, rather than excluding them.
@@ 3.2.6 Agent Interaction Pattern
While prior works have typically reported on dataset-level statistics (e.g., basic information in Table 1), it has rarely addressed how agents actually interact within their environments. We additionally report the following information, providing an overview of the interactive tendency of employed agents in their specific scenarios.
The Number of Turns:
Different from offline dialogue datasets where a deterministic number of turns is provided, recently employed online evaluation scenarios typically omit details about the agent interactivity. However, the number of interactive turns serves as a critical indicator of the conversational or interactive dynamics. In general, planning scenarios with more turns tend to require further management of user constraints, suggesting the need for sustained engagement and state tracking.
The Number of Steps:
Steps refer to internal actions or reasoning hops, such as goal decomposition or tool use. This feature reflects the agent’s ability to handle complex tasks that require breaking down overarching goals into smaller, manageable actions. Fewer steps may suggest shallow or generic reasoning, while more steps indicate finer control. Correlating the count of steps with performance and the length of interaction can reveal whether agents are making meaningful progress or getting stuck in unproductive loops.
@@ 4.1 Benchmarks
To verify our hypothesis on the evaluation following behavioral stages, we employ challenging task planning benchmarks and compare agent performance based solely on task completion with that achieved using AURA. Specifically, we utilize three task domains of two benchmarks: TravelPlanner (pmlr-v235-xie24j), a single-turn dataset in the itinerary domain, and $\tau$-Bench (yao2025taubench), a multi-turn dataset encompassing airline and retail domains. Table 1 presents the statistics of these benchmarks. These benchmarks are selected due to their rich environmental constraints that conform to the prerequisite assumptions described in Section 3 (i.e., predefined tools and external database). For the evaluation metric for task completion ($\mathcal{R}$), pass rate is used for TravelPlanner, and passˆk is used in $\tau$-Bench. Refer to Table 5 for the description of each metric.
4.2 LLMs for Agent, User Simulator, and Task Evaluator
Table 2: AURA evaluation result. The relative performance ranking among agents for each metric is differentiated with colors (i.e., darker color indicates more competitive performance; different colors are applied for proprietary and open-source models). Best average performance is indicated in bold, and second-best is underlined.
@@ 4.2 LLMs for Agent, User Simulator, and Task Evaluator
Agent $\mathcal{S}$ $\mathcal{A}$ $\mathcal{O}$ $\mathcal{P}$ $\mathcal{R}$ AVG. $\mathcal{S}$ $\mathcal{A}$ $\mathcal{O}$ $\mathcal{P}$ $\mathcal{R}$ AVG. $\mathcal{S}$ $\mathcal{A}$ $\mathcal{O}$ $\mathcal{P}$ $\mathcal{R}$ AVG.
@@
We test diverse task planning agents consisting of five proprietary models (gpt-4o, gpt-4o-mini, gpt-3.5-turbo, gemini-1.5-flash, sonnet-3.5) and four open-weight models (mistral-large 222123B-sized model: https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, mixtral-8x7B, llama-3.3-70B, qwen2.5-72B), with a temperature of 0.0, respectively. Since $\tau$-Bench requires multi-turn interactions between a user and agent, we employ gpt-4o as a user simulator with the same instruction as the original work (yao2025taubench). For the LLM evaluator leveraged in measuring AURA metrics, we employ llama-3.3-70B (grattafiori2024llama3herdmodels) [^3] We confirm llama-3.3-70B shows better performance compared to gpt-4o-mini through our preliminary investigation.. All prompts used in our experiments are listed in Appendix B.
@@
Agent Avg. Turns Avg. Steps Avg. Turns Avg. Steps Avg. Turns Avg. Steps
@@
Open-Weight Models
@@ 4.3 Results
Behavioral Stage Diagnosis:
Table 2 presents the evaluation results following AURA (see Appendix C for agents’ erroneous behaviors at each stage). We differentiate the color ranking of proprietary and open-weight models to visually observe patterns. Along with individual metrics in AURA, we report the average score to see the overall behavioral performance.
A notable observation is that the conventional task completion metric ($\mathcal{R}$) does not necessarily reflect the performance of intermediate phases. For instance, while qwen2.5-72B achieves the highest performance on task completion (.38) among open-weight models in $\tau$-Bench-Retail, another AURA results lag behind other models (e.g., llama-3.3-70B). Moreover, although most agents in TravelPlanner show the same zero task completion performance [^4] This low task completion result is consistent with the original paper’s results (pmlr-v235-xie24j)., each model has different intermediate capability patterns according to AURA metrics. These observed patterns imply a potential discrepancy between task completion performance values and qualitative user preferences, as user satisfaction is influenced not only by task completion but also by the quality of interactions across all phases. We introduce a human study for verifying this discrepancy in Section 5.1.
Interaction Pattern Diagnosis:
We summarize each agent’s interaction pattern in Table 3. While TravelPlanner is characterized by single-turn interactions with numerous intermediate steps prior to producing a final response, $\tau$-Bench exhibits fewer intermediate steps per turn but involves multiple turns. These patterns reveal certain behavioral tendencies among the agents; for instance, competitive agents (e.g., sonnet-3.5, llama-3.3-70B) appear to engage in more extensive internal reasoning within individual turns, yet participate in fewer turns overall in multi-turn benchmarks. This observation provides insightful suggestions on scenario diversification. For example, when adapting TravelPlanner for multi-turn scenarios, careful consideration must be given to balancing the number of turns and steps within each turn. Given that service providers often target different application scenarios, analyzing interaction patterns across benchmarks can inform the strategic selection of those most aligned with specific deployment contexts.
@@ 5 Analyses and Discussions
5.1 Relationship Between Task Completion, AURA, and User Satisfaction
Our hypothesis regarding the results presented in Table 2 is that factors beyond task completion may have significantly influenced overall user satisfaction. To investigate this through a human study, we recruit 16 graduate-level participants, a sample size congruent to prior human-computer interaction research 10.1145/1735223.1735255; 10.1145/2858036.2858498. This study is conducted using two controlled scenarios, described as follows:
(1) Same $\mathcal{R}$ & Different AVG.
We select gpt-4o-mini and mistral-large in $\tau$-Bench-Airline, as these two models exhibit identical task completion performance (.26), yet differ in their AURA average scores (.64 and .56, respectively).
@@ 5.1 Relationship Between Task Completion, AURA, and User Satisfaction
We select gemini-1.5-fsh. and llama-3.3-70B in $\tau$-Bench-Retail, as these two models exhibit identical AURA average score (.64), yet differ in their task completion performance (.15 and .36, respectively).
Figure 3: Human study results examining preferred interactions.
Participants are presented with two multi-turn conversations generated by different agents interacting with a gpt-4o user simulator. They are asked to indicate which conversation they find more satisfactory. If both conversations are perceived as equally (un)satisfactory, participants are given the option to select “no preference.” Each participant evaluates a total of six conversation pairs (randomly sampled three pairs of conversations, respectively). Detailed instructions are demonstrated in Section D.1, and we provide qualitative analysis for participants’ decisions in Section D.2.
$\mathcal{S}$ $\mathcal{A}$ $\mathcal{O}$ $\mathcal{P}$ $\mathcal{R}$ AVG.
@@
The results of the human evaluation are presented in Figure 3. For each scenario, we conduct binomial tests by excluding “no preference” responses, yielding statistically significant results (p-value << 0.05 in both cases). In study (1), we observe that participants tend to favor conversations with higher average score when task completion is held constant. Conversely, in study (2), task completion still remains a strong predictor of user satisfaction when the average score is identical. These findings highlight the complementary roles of general task completion and AURA metrics in capturing different facets of user satisfaction, as demonstrated in Table 2. Our study thus supports the use of AURA as a meaningful proxy for user satisfaction and motivates future research into evaluation frameworks that go beyond simple task outcomes.
5.2 Can Mixing Agents Lead to Better Performance?
Building upon the insights from previous experiments, we are further motivated to explore the potential for enhancing performance and user satisfaction by leveraging diverse agents strengths within the agentic pipeline. To facilitate this investigation, we separate the agent’s interactive process into two distinct components: (1) intermediate understanding and (2) response generation. This separation enables targeted improvements, wherein state consistency and tool efficiency are addressed in the first stage, whereas observation alignment and policy adherence are the focus of the second. Specifically, we employ one agent to understand user requests and call proper tools, then replace it with the other when starting to generate responses.
For faster investigation of this experimental study, we run FP8-quantized versions of three open-weight models (llama-3.3-70B, qwen2.5-72B, mistral-large), alongside llama-3.3-70B as the user simulator, and evaluate on $\tau$-Bench-Airline with five repeated runs per configuration. Selected results are presented in Table 4. While acknowledging that arbitrary combinations of agents do not consistently yield improved performance, even leading to drop in performance, certain pairings demonstrate notable enhancements in either task completion or the average AURA score. Notably, combining qwen2.5-72B (for intermediate understanding) with llama-3.3-70B (for response generation) leads to a substantial improvement in task completion (.263 $\rightarrow$ .361) without compromising the average AURA score too much (.628 $\rightarrow$ .624). These findings suggest a promising direction for future research to strategically arrange agents across different stages of the task planning pipeline. For additional details and comprehensive experimental results beyond the scope of this discussion, please refer to Appendix E.
@@ 6 Conclusion
We propose AURA, a task-agnostic evaluation framework based on a partially observable Markov Decision Process (POMDP), to more accurately capture how user satisfaction emerges from both intermediate behaviors and final outcomes in interactive task planning. Unlike traditional metrics that focus solely on task completion, AURA provides a holistic lens for assessing agent performance by modeling the user’s evolving experience throughout the interaction. Our analysis reveals that different models excel at different stages of a task, underscoring the limitations of evaluating performance through a single end-point measure. Human studies further demonstrate that stage-specific behavioral indicators often correlate more strongly with overall user satisfaction than task completion alone. These findings highlight the importance of comprehensive evaluation methods and establish AURA as a principled, diagnostic tool for developing more transparent, robust, and user-aligned planning agents.
Limitations and Future Directions
While we introduce AURA as a diagnostic framework for estimating user satisfaction, we acknowledge its limitations that offer avenues for future research. First, AURA relies on LLMs to serve as the evaluator for its atomic criteria. Although the assessment employs a model based on boolean criteria, demonstrated to be more reliable than long-text evaluations, the accuracy and consistency of the results may nonetheless depend on the specific capabilities and potential bias of the selected model. Furthermore, as discussed in Section 5.3, ensuring the reliability and realism of user simulators remains an open challenge, warranting further investigation. As a future direction, we believe that AURA can provide fine-grained rewards to enhance human-LLM collaboration, thereby further improving user satisfaction and demonstrating synergies with collaborative simulation wu2025collabllm.
@@ Appendix A Evaluation Metrics Across Task Planning Benchmarks
MultiWOZ [budzianowski-etal-2018-multiwoz] Success Rate The system answered all requested attributes.
@@
AgentBench [liu2024agentbench] Domain-Specific Different metrics such as Success Rate, Win Rate,
@@
WebLINX [10.5555/3692070.3693410] Intent Match Boolean indicator for correct intent prediction.
@@
FlowBench [xiao-etal-2024-flowbench] Tool Invocation Correctly identified tool configs (P/R/F1).
@@
$\tau$-Bench [yao2025taubench] Pass@k Boolean success indicator for k attempts.
Passˆk Whether all k trials are successful.
@@
WebShop [NEURIPS2022_82ad13ec] Task Score The average reward across episodes.
@@
TravelPlanner [pmlr-v235-xie24j] Delivery Rate The agent completed tasks within limited steps.
@@ Appendix B Prompts
Basically, we keep the same agent prompt as the original paper [pmlr-v235-xie24j, yao2025taubench]. The following prompts are evaluation prompts adopted in AURA evaluation.
@@ 3. Accuracy and truthfulness
[ { "state_id": "1", "justification": "brief explanation...", "score": "0" }, { "state_id": "2", "justification": "brief explanation...", "score": "1" } ]
@@ 4. Consistency with previous details
[ { "response_id": "1", "justification": "brief explanation...", "score": "0" }, { "response_id": "2", "justification": "brief explanation...", "score": "1" } ]
@@
[ { "policy_id": "1", "justification": "Mention specific policy and turn message, if violated. Otherwise, mention all policies are adhered to.", "score": "0" }, { "policy_id": "2", "justification": "Mention specific policy and turn message, if violated. Otherwise, mention all policies are adhered to.", "score": "1" } ]